# How to build the best medical image segmentation algorithm using foundation models: a comprehensive empirical study with Segment Anything Model

### v2

@@ 4.2.1 Implementation details
When implementing task-expansive prompt-based learning, we utilized a collection of datasets and their corresponding masks. Our experimental setup focused on training with all MRI datasets listed in Table 2 except “MRI-Knee”. In this collection criteria, we include 13 MRI datasets including 45k training image and non-empty mask pairs, and 7.9k evaluation images and mask pairs. We selected ViT-B as the encoder architecture since (1) it achieved similar performance to ViT-H in the data-specific fine-tuning setting; (2) MobileSAM was a distilled version of SAM which was not provided by initial SAM and thus less representative.
Under this multi-dataset supervised training setting, we configured a batch size of 3, utilizing DDP across 3 A6000 GPUs for training. The training was stopped when the validation performance was not improved for 5 epochs. To train the network, we fixed the prompt encoder and updated the image encoder and mask decoder in a vanilla setting (same as SAM’s supervised training stage). We generated a box or point prompt for each image as described in Section 2.3. Additionally, in this setting, we incorporated MedSAM [59] as a parallel approach since its training algorithm was similar to the proposed task-expansive prompt-based learning but with a different database.
In this experimental setup, our methodology follows a two-phase procedure, depicted in Figure 1 Setup 2: (1) Phase 1: task-expansive prompt-based learning. (2) Phase 2: task-specific supervised training for each dataset, beginning from pre-trained weights. We name the weights after phase 1 as SupSAM. The implementation details in the task-specific setting were identical to those of the previous section (Section 4.1). In phase 2, we had 2 model component choices (En/Decoder or Decoder only) and 3 Vanilla/PEFT choices in the automated setting. We evaluated the final automated segmentation performance after phase 2.
@@ 4.2.2 Results
When fine-tuning both the encoder and decoder during phase 2 task-specific supervised training (Figure 6 and 7, bottom 3 blocks), we observe either a negative or no effect when utilizing a different pre-trained weights when using PEFT methods and a positive effect when using Vanilla fine-tuning. However, the latter setting leads to worse overall performance and thus offers little practical value. Despite the decrement being alleviated when using MedSAM, the overall trend shows the ineffectiveness of prompt-based learning when the ultimate goal is automated segmentation (MedSAM has around 34 times larger training set yet results in no performance change at best). We infer that the observed performance enhancement in En/Decoder-Vanilla is attributed to supervised pre-training potentially nudging the weights closer to the medical domain, thereby reducing the likelihood of model collapse. Nonetheless, given that this configuration still results in the lowest segmentation performance compared to the PEFT setting, this improvement does not offer practical benefit.
Our research indicates that supervised pre-training across various labeled medical images, despite working well in the original setting (MedSAM [59] is verified in its paper and ours is verified in Section 4.2), fails to enhance automated segmentation performance. We believe the ineffectiveness of additional supervised pre-training is due to the difference in the tasks: (1) The weights learned from interactive (prompt-based) mode vs. automated mode might be different and could not commute with each other effectively. Additionally, extended pre-training might disrupt the original weights’ transferability. (2) Furthermore, training the network on multiple segmentation tasks simultaneously could lead to conflicts within the training process thus making the training more challenging.
@@ 4.3.1 Implementation details
In this stage, all MRI datasets in Table 2 were used during training. The mask information was ignored and only the images from the training set were utilized. Blank slices were further removed since they were not informative for the reconstruction task. This resulted in images for training.
Following the previous section, we only select ViT-B as the encoder architecture. To employ MAE’s objective, we introduced an additional decoder to reconstruct the inputs, which was discarded after training. The additional decoder’s structure, mask ratio, and learning rate schedule all follow that of the original paper. Since the shape of SAM’s input was restricted to , we set the batch size to . To accommodate for the small batch size, we set the accumulative iteration, i.e. the number of iterations to only compute gradients without updating the network parameters, to . The training was conducted on 8 A6000 for epochs.
@@ 4.3.2 Results
Figure 8 presents the results of using SSLSAM as the weights before task-specific supervised training. We observe an improvement in the average performance of all MRIs across all architecture combinations besides En/Decoder + LoRA. Noticeably, when selecting En/Decoder + Adapter as the fine-tuning strategy, we observe consistent improvements over all MRI datasets, and the model achieves an average performance of , which surpasses the leading ViT-H based method by and the best UNet-based method (nnUNet) by . Contradictory to task-expansive prompt-based learning, the features learned during SSL are task agnostic, and since we only pre-train the encoder, there is no dependence on the task condition, which we have shown to hurt the performance. In the non-MRI data sets base, we observe a minor decrease in performance. This may be due to the specificity of the MRI.
@@ 4.4.1 Implementation details
Our study primarily investigates 5-shot learning, selecting five images and their corresponding annotations randomly from the training set and validation set, respectively. In the case of datasets derived from 3D volumes, these samples were chosen from distinct volumes to ensure diversity. In addition to the constraint of using only 5 training/validation images, we also modified the training hyper-parameters better suited for the few-shot setting. Specifically, we set a maximum of 1000 epochs for training, a batch size of 2, and did not include any early stopping criteria.
In this setting, we also investigated the previous three setups within a few-shot framework, including (1) task-specific fine-tuning exclusively, and (2) two pre-trained setups: task-expansive prompt-based learning and task-agnostic self-supervised learning, each followed by additional task-specific fine-tuning. Notably, we cannot use the weights learned from task-expansive prompt-based learning as its training pipeline requires all labeled images. Thus, we consider the MedSAM weights as an alternative. In this section, we investigate three key questions: (1) whether additional pre-training offers superior initial weights for the few-shot setting, (2) which fine-tuning strategy proves most effective in a few-shot context, and (3) whether fine-tuning foundation models still work better than UNets in the few-shot setting?
@@ 4.4.2 Results
Figure 9 shows the results for 5-shot setting, comparing the fine-tuning of SAM’s models with direct training on UNets. The results demonstrate that using ViT-B + En/Decoder + Adapter on the initial SAM yields optimal performance across all configurations. Comparing the results with UNets’, especially with the best-performing nnUNet, fine-tuned SAM still manages to achieve a slight improvement. This observation aligns with the findings presented in Section 4.1.
Consistently, the observation that MedSAM is ineffective persists in this few-shot scenario. While SSLSAM has shown effectiveness in the full-dataset setting, its efficacy diminished and even became detrimental in the few-shot scenario. This observation aligns with the understanding that the task-agnostic training disrupts the model’s connection with the segmentation tasks. While beneficial as initial weights for full-dataset training, this approach may not adequately rebuild these connections with only five images, resulting in lower performance. In summary, we observe the ineffectiveness of additional pre-training (either using task-expansive prompt-based learning or task-agnostic self-supervised learning) with limited training data. This can be contradictory to the common belief that imposing medical knowledge to the network in advance will be beneficial in the few-shot learning scenario.
@@ 4.5 What if I can provide prompts to SAM?
Despite the primary goal of this paper is to find the best practice to build an automated segmentation algorithm from the foundation models, we conduct a non-trivial extension that studies the effectiveness of various task-specific supervised training strategies in the interactive segmentation setting.
@@ 4.5.1 Implementation details
We included the performance of applying prompts to the original SAM and its oracle performance; the latter reflects the hypothetical performance when conducting an extra step to determine the best prediction from the network. Additionally, we include weights acquired from task-expansive prompt-based learning and MedSAM in the box-prompt setting. When applying task-specific supervised training to SAM, we followed Section 4.1 to include all strategies.
@@ 4.5.2 Results
Figure 10 shows the performance of all task-specific supervised training strategies in the interactive mode, as well as that of the original SAM as a reference. First, we observe the performance of SAM with point prompts is not satisfying. Replacing the point prompts with box ones significantly improves the performance. In this setting, we also examine the weights acquired from task-expansive prompt-based learning and observe similar performance (less than DSC difference compared to that of the initial SAM). This observation confirms that the inefficiency of task-expansive prompt-based learning is inherently from having different tasks in mind.
@@
Conducting task-specific supervised training is also shown to be effective in the interactive segmentation scenario, especially when points are used as the prompts. Moreover, we find the best task-specific supervised training strategy is ViT-T + En/Decoder + Adapter in both scenarios, achieving DSC when using a point/box as the prompt.
@@ 5 Conclusion
In summary, we systematically study different strategies to fine-tune the Segment Anything Model (SAM) in the medical imaging field. Our extensive experiments reveal several interesting observations and provide a guideline for developing an automated segmentation algorithm from SAM. Our findings are based on dataset availability, and are summarized as follows:
@@
Fine-tuning SAM, if with proper strategies, leads to better segmentation performance than training from scratch with nnUNet [40].
@@
Increasing the size of network architecture leads to a marginal increase in segmentation performance while exponentially longer training time.
@@
Using parameter-efficient fine-tuning (PEFT) techniques has a clear advantage over fine-tuning all parameters on medium-to-large-size architectures (ViT-H or ViT-B) but not on the small architecture (ViT-T).
@@
Using additional unlabeled data with self-supervised learning can improve SAM’s performance when the test data is from the same modality as the pre-trained data and has no effect otherwise.

### v3

@@ 4.2.1 Implementation details
When implementing task-expansive prompt-based learning, we utilized a collection of datasets and their corresponding masks. Our experimental setup focused on training with all MRI datasets listed in Table 2 except “MRI-Knee”. In this collection criteria, we include 13 MRI datasets, including 45k training images and non-empty mask pairs, and 7.9k evaluation images and mask pairs. We selected ViT-B as the encoder architecture since (1) it achieved similar performance to ViT-H in the data-specific fine-tuning setting; (2) MobileSAM was a distilled version of SAM, which was not provided by the initial SAM and thus less representative.
Under this multi-dataset supervised training setting, we configured a batch size of 3, utilizing DDP across 3 A6000 GPUs for training. The training was stopped when the validation performance did not improve for 5 epochs. To train the network, we fixed the prompt encoder and updated the image encoder and mask decoder in a vanilla setting (same as SAM’s supervised training stage). We generated a box or point prompt for each image as described in Section 2.3. Additionally, in this setting, we incorporated MedSAM (Ma et al., 2024) as a parallel approach since its training algorithm was similar to the proposed task-expansive prompt-based learning but with a different database.
In this experimental setup, our methodology follows a two-phase procedure, depicted in Figure 1. Setup 2: (1) Phase 1: task-expansive prompt-based learning. (2) Phase 2: task-specific supervised training for each dataset, beginning from pre-trained weights. We name the weights after phase 1 as SupSAM. The implementation details in the task-specific setting were identical to those of the previous section (Section 4.1). In phase 2, we had 2 model component choices (En/Decoder or Decoder only) and 3 Vanilla/PEFT choices in the automated setting. We evaluated the final automated segmentation performance after phase 2.
@@ 4.2.2 Results
When fine-tuning both the encoder and decoder during phase 2 task-specific supervised training (Figure 6 and 7, bottom 3 blocks), we observe either a negative or no effect when utilizing a different pre-trained weights when using PEFT methods and a positive effect when using Vanilla fine-tuning. However, the latter setting leads to worse overall performance and thus offers little practical value. Despite the decrement being alleviated when using MedSAM, the overall trend shows the ineffectiveness of prompt-based learning when the ultimate goal is automated segmentation, as MedSAM has a training set around 34 times larger yet results in no performance change at best. We infer that the observed performance enhancement in En/Decoder-Vanilla is attributed to supervised pre-training, potentially nudging the weights closer to the medical domain, thereby reducing the likelihood of model collapse. Nonetheless, since this configuration still yields the lowest segmentation performance compared to the PEFT setting, this improvement does not offer a practical benefit.
Our research indicates that supervised pre-training across various labeled medical images, despite working well in the original setting (MedSAM (Ma et al., 2024) is verified in its paper, and ours is verified in Section 4.2), fails to enhance automated segmentation performance. We believe the ineffectiveness of additional supervised pre-training is due to the difference in the tasks: (1) The weights learned from interactive (prompt-based) mode vs. automated mode might be different and could not communicate with each other effectively. Additionally, extended pre-training may disrupt the transferability of the original weights. (2) Furthermore, training the network on multiple segmentation tasks simultaneously could lead to conflicts within the training process, thus making the training more challenging. These observations suggest that hybrid training, i.e., fine-tuning SAM for both the automatic segmentation task and the prompt-based segmentation task simultaneously, may not be effective.
@@ 4.3.1 Implementation details
In this stage, all MRI datasets in Table 2 were used during training. The mask information was ignored, and only the images from the training set were utilized. Blank slices were further removed since they were not informative for the reconstruction task. This resulted in images for training.
Following the previous section, we only select ViT-B as the encoder architecture. To employ MAE’s objective, we introduced an additional decoder to reconstruct the inputs, which was discarded after training. The additional decoder’s structure, mask ratio, and learning rate schedule all follow that of the original paper. Since the shape of SAM’s input was restricted to , we set the batch size to . To accommodate for the small batch size, we set the cumulative iteration, i.e. the number of iterations to only compute gradients without updating the network parameters, to . The training was conducted on 8 A6000 for epochs.
@@ 4.3.2 Results
Figure 8 presents the results of using SSLSAM as the weights before task-specific supervised training. We observe an improvement in the average performance of all MRIs across all architecture combinations besides En/Decoder + LoRA. Noticeably, when selecting En/Decoder + Adapter as the fine-tuning strategy, we observe consistent improvements over all MRI datasets, and the model achieves an average performance of , which surpasses the leading ViT-H based method by and the best UNet-based method (nnUNet) by . Contradictory to task-expansive prompt-based learning, the features learned during SSL are task agnostic, and since we only pre-train the encoder, there is no dependence on the task condition, which we have shown to hurt the performance. In the non-MRI data sets base, we observe a minor decrease in performance. This may be due to the MRI’s specificity.
@@ 4.4.1 Implementation details
Our study primarily investigates 5-shot learning, where we randomly select five images and their corresponding annotations from the training set and validation set, respectively. In the case of datasets derived from 3D volumes, these samples were chosen from distinct volumes to ensure diversity. In addition to the constraint of using only 5 training and validation images, we also modified the training hyperparameters to suit the few-shot setting better. Specifically, we set a maximum of 1000 epochs for training and a batch size of 2, and did not include any early stopping criteria. We increased the number of epochs to ensure sufficient updates for learning medical image patterns, used a smaller batch size to have diverse updates, and removed early stopping to prevent premature termination in the few-shot setting. We chose this strategy based on preliminary results indicating better performance compared to the default setting for full-set fine-tuning.
In this setting, we also investigated the previous three setups within a few-shot framework, including (1) task-specific fine-tuning exclusively and (2) two pre-trained setups: task-expansive prompt-based learning and task-agnostic self-supervised learning, each followed by additional task-specific fine-tuning. Notably, we cannot use the weights learned from task-expansive prompt-based learning as its training pipeline requires all labeled images. Thus, we consider the MedSAM weights as an alternative. In this section, we investigate three key questions: (1) whether additional pre-training offers superior initial weights for the few-shot setting, (2) which fine-tuning strategy proves most effective in a few-shot context, (3) whether fine-tuning foundation models still works better than UNets in the few-shot setting, and (4) how does performance vary with different levels of few-shot data availability, such as 10-shot or 20-shot?
@@ 4.4.2 Results
Figure 9 shows the results for the 5-shot setting, comparing the fine-tuning of SAM’s models with direct training on UNets. The results demonstrate that using ViT-B + En/Decoder + Adapter on the initial SAM yields optimal performance across all configurations. Comparing the results with UNets’, especially with the best-performing nnUNet, fine-tuned SAM still manages to achieve a slight improvement. This observation aligns with the findings presented in Section 4.1. The performance of three selected datasets with different levels of few-shot data availability is shown in Appendix Figure 12.
Consistently, the observation that MedSAM is ineffective persists in this few-shot scenario. While SSLSAM has shown effectiveness in the full-dataset setting, its efficacy diminished and even became detrimental in the few-shot scenario. This observation aligns with the understanding that the task-agnostic training disrupts the model’s connection with the segmentation tasks. While beneficial as initial weights for full-dataset training, this approach may not adequately rebuild these connections with only five images, resulting in lower performance. In summary, we observe the ineffectiveness of additional pre-training (either using task-expansive prompt-based learning or task-agnostic self-supervised learning) with limited training data. This can be contradictory to the common belief that imposing medical knowledge on the network in advance will be beneficial in the few-shot learning scenario.
| Dataset Setting | Condition | Optimal Strategy |
| Single Dataset | NA | ViT-H + En/Decoder + LoRA |
| Multiple Datasets | NA | SSLSAM + En/Decoder + Adapter |
| SIngle Dataset | Few-shot | VIT-B + En/Decoder + Adapter |
| Single Dataset | Prompt-based | ViT-T + En/Decoder + Adapter |
@@ 4.5 What if I can provide prompts to SAM?
Despite the primary goal of this paper being to find the best practice to build an automated segmentation algorithm from the foundation models, we conduct a non-trivial extension that studies the effectiveness of various task-specific supervised training strategies in the interactive segmentation setting.
@@ 4.5.1 Implementation details
We included the performance of applying prompts to the original SAM and its oracle performance; the latter reflects the hypothetical performance when an extra step is taken to determine the best prediction from the network. Additionally, we include weights acquired from task-expansive prompt-based learning and MedSAM in the box-prompt setting. When applying task-specific supervised training to SAM, we followed Section 4.1 to include all strategies.
@@ 4.5.2 Results
Figure 10 shows the performance of all task-specific supervised training strategies in the interactive mode, as well as that of the original SAM as a reference. First, we observe that the performance of SAM with point prompts is not satisfactory. Replacing the points with boxes significantly improves performance. In this setting, we also examine the weights acquired from task-expansive prompt-based learning and observe a similar performance, with a difference of only in DSC compared to the initial SAM. This observation confirms that the inefficiency of task-expansive prompt-based learning is inherently due to having different tasks in mind.
@@
Conducting task-specific supervised training is also shown to be effective in the interactive segmentation scenario, especially when points are used as the prompts. Moreover, we find that the best task-specific supervised training strategy is ViT-T + En/Decoder + Adapter in both scenarios, achieving DSC when using a point/box as a prompt. This difference also suggests that providing box prompts can lead to better segmentation performance with the additional annotation cost.
@@ 5 Conclusion
In summary, we systematically study different strategies to fine-tune the Segment Anything Model (SAM) in the medical imaging field. Our extensive experiments reveal several interesting observations and provide a guide for developing an automated segmentation algorithm from SAM. Our findings are based on dataset availability and are summarized as follows:
@@
Fine-tuning SAM, if with proper strategies, leads to better segmentation performance than training from scratch with nnUNet (Isensee et al., 2021).
@@
Increasing the size of network architecture leads to a marginal increase in segmentation performance, while exponentially longer training time.
@@
Using parameter-efficient fine-tuning (PEFT) techniques has a clear advantage over fine-tuning all parameters on medium-to-large-sized architectures (ViT-H or ViT-B), but not on small architectures (ViT-T).
@@
Using additional unlabeled data with self-supervised learning can improve SAM’s performance when the test data is from the same modality as the pre-trained data, and has no effect otherwise.
@@
In addition to these findings, we also observe several limitations when fine-tuning SAM.
Namely, the computational cost of fine-tuning SAM can be high when using the default backbones. As shown in Table 3, when using ViT-H or ViT-B as the backbone, the number of parameters is 640.95M and 93.60 M. Using parameter-efficient fine-tuning strategies leads to better segmentation performance but does not improve the training efficiency. A potential solution is to utilize smaller backbones, such as ViT-T, i.e., a lightweight encoder distilled from pre-trained ViT-H (Zhang and Jiao, 2023), to balance the performance and computation cost.
Another limitation is that our study only focuses on the 2D segmentation tasks.
To alleviate this issue, we have conducted a preliminary study that utilizes nnUNet 3D as a strong 3D segmentation baseline. We observed that the performance of DSC(%) is 87.14 vs. 87.35, 40.56 vs. 29.96, 47.58 vs. 46.95 on MRI-Heart, MRI-Scler., and CT-Colon, respectively, when comparing between nnUNet 3D vs. 2D. This observation suggests that the improvements are limited when the baseline methods are switched from 2D to 3D.
On the other hand, adapting SAM to 3D segmentation is beyond the scope of this paper. We believe some potential directions include adapting SAM directly to 3D (Gong et al., 2024) or fine-tuning the recently proposed SAM 2 (Ravi et al., 2024), which is designed specifically for video segmentation.
Research reported in this publication was supported by the National Institute Of Biomedical Imaging And Bioengineering of the National Institutes of Health under Award Number R01EB031575. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
The conflicts of interest have not been entered yet.
All data used in this study are publicly available. We provide detailed descriptions and proper citations for each dataset used to ensure full transparency and reproducibility. Additionally, all code for data preprocessing and preparation is publicly released and can be used to fully reproduce our results. The code is available at: https://github.com/mazurowski-lab/finetune-SAM.
Appendix A Proof of the central theorem
In this appendix we prove the central theorem and present additional experimental results.
Remainder omitted in this sample.
Appendix B Appendix section
In the appendix, we provide the numerical results of all experiments.
Tables 5, 6, 7 present the results of applying task-specific supervised training to ViT-H, ViT-B, and ViT-T respectively, and table 8 shows the performance of UNets. Figure 4 summarizes the results of these 4 tables, and Figure 11 visualizes some examples of segmentation results of using foundation model vs. using nnUNet.
Tables 9, 10, 11 show the performance of replacing SAM’s weights with SupSAM (acquired from task-expansive prompt-based learning), MedSAM (from Ma et al. (2024)), and SSLSAM (acquired from task-agnostic self-supervised learning), respectively. Since all weights utilize the same encoder architecture (ViT-B), we show the difference by conducting Table 6 minus the corresponding table in Figure 6, 7, and 8, respectively.
Tables 13 and 12 show the performance of applying task-specific supervised training in the few-shot learning setting, which is summarized in Figure 9. The results for the selected three datasets of expanded few-shot learning: 10-shot, 20-shot, are shown in Figure 12.
Table 14 shows the performance of initial SAM in the point-based segmentation setting, with tables 15, 16, and 17 showing applying task-specific supervised training in the same setting. Tables 18, 19, 20, and 21 are conducted in the same setting but with the box as the prompt. All results are summarized in Figure 10.
| Dataset | ViT-H | |||||||||||||
| Decoder | En/Decoder | |||||||||||||
| V | A | L | V | A | A* | L | ||||||||
| DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | |
| MRI-Prost. | ||||||||||||||
| MRI-Heart | ||||||||||||||
| MRI-Brain | ||||||||||||||
| MRI-Scler. | ||||||||||||||
| MRI-Hippo. | ||||||||||||||
| MRI-Kidney | ||||||||||||||
| MRI-Stroke. | ||||||||||||||
| MRI-Resect. | ||||||||||||||
| MRI-Breast | ||||||||||||||
| CT-Colon | ||||||||||||||
| CT-Panc. | ||||||||||||||
| CT-vessel. | ||||||||||||||
| US-Breast | ||||||||||||||
| US-Muscle | ||||||||||||||
| US-Kidney | ||||||||||||||
| Xray-Hip | ||||||||||||||
| Xray-Chest | ||||||||||||||
| Average | ||||||||||||||
| Dataset | ViT-B | |||||||||||
| Decoder | En/Decoder | |||||||||||
| V | A | L | V | A | L | |||||||
| DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | |
| MRI-Prost. | ||||||||||||
| MRI-Heart | ||||||||||||
| MRI-Brain | ||||||||||||
| MRI-Scler. | ||||||||||||
| MRI-Hippo. | ||||||||||||
| MRI-Kidney | ||||||||||||
| MRI-Stroke. | ||||||||||||
| MRI-Resect. | ||||||||||||
| MRI-Breast | ||||||||||||
| CT-Colon | ||||||||||||
| CT-Panc. | ||||||||||||
| CT-vessel. | ||||||||||||
| US-Breast | ||||||||||||
| US-Muscle | ||||||||||||
| US-Kidney | ||||||||||||
| Xray-Hip | ||||||||||||
| Xray-Chest | ||||||||||||
| Average | ||||||||||||
| Dataset | ViT-T | |||||||||||
| Decoder | En/Decoder | |||||||||||
| V | A | L | V | A | L | |||||||
| DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | |
| MRI-Prost. | ||||||||||||
| MRI-Heart | ||||||||||||
| MRI-Brain | ||||||||||||
| MRI-Scler. | ||||||||||||
| MRI-Hippo. | ||||||||||||
| MRI-Kidney | ||||||||||||
| MRI-Stroke. | ||||||||||||
| MRI-Resect. | ||||||||||||
| MRI-Breast | ||||||||||||
| CT-Colon | ||||||||||||
| CT-Panc. | ||||||||||||
| CT-vessel. | ||||||||||||
| US-Breast | ||||||||||||
| US-Muscle | ||||||||||||
| US-Kidney | ||||||||||||
| Xray-Hip | ||||||||||||
| Xray-Chest | ||||||||||||
| Average | ||||||||||||
| Dataset | UNet | nnUNet | SwinUNet | |||
| DSC | NSD | DSC | NSD | DSC | NSD | |
| MRI-Prost. | ||||||
| MRI-Heart | ||||||
| MRI-Brain | ||||||
| MRI-Scler. | ||||||
| MRI-Hippo. | ||||||
| MRI-Kidney | ||||||
| MRI-Stroke. | ||||||
| MRI-Resect. | ||||||
| MRI-Breast | ||||||
| CT-Colon | ||||||
| CT-Panc. | ||||||
| CT-vessel. | ||||||
| US-Breast | ||||||
| US-Muscle | ||||||
| US-Kidney | ||||||
| Xray-Hip | ||||||
| Xray-Chest | ||||||
| Average | ||||||
| Dataset | SupSAM | |||||||||||
| Decoder | En/Decoder | |||||||||||
| V | A | L | V | A | L | |||||||
| DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | |
| MRI-Prost. | ||||||||||||
| MRI-Heart | ||||||||||||
| MRI-Brain | ||||||||||||
| MRI-Scler. | ||||||||||||
| MRI-Hippo. | ||||||||||||
| MRI-Kidney | ||||||||||||
| MRI-Stroke. | ||||||||||||
| MRI-Resect. | ||||||||||||
| MRI-Breast | ||||||||||||
| CT-Colon | ||||||||||||
| CT-Panc. | ||||||||||||
| CT-Vessel | ||||||||||||
| US-Breast | ||||||||||||
| US-Muscle | ||||||||||||
| US-Kidney | ||||||||||||
| Xray-Hip | ||||||||||||
| Xray-Chest | ||||||||||||
| Average | ||||||||||||
| Dataset | MedSAM | |||||||||||
| Decoder | En/Decoder | |||||||||||
| V | A | L | V | A | L | |||||||
| DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | |
| MRI-Prost. | ||||||||||||
| MRI-Heart | ||||||||||||
| MRI-Brain | ||||||||||||
| MRI-Scler. | ||||||||||||
| MRI-Hippo. | ||||||||||||
| MRI-Kidney | ||||||||||||
| MRI-Stroke. | ||||||||||||
| MRI-Resect. | ||||||||||||
| MRI-Breast | ||||||||||||
| CT-Colon | ||||||||||||
| CT-Panc. | ||||||||||||
| CT-Vessel | ||||||||||||
| US-Breast | ||||||||||||
| US-Muscle | ||||||||||||
| US-Kidney | ||||||||||||
| Xray-Hip | ||||||||||||
| Xray-Chest | ||||||||||||
| Average | ||||||||||||
| Dataset | SSLSAM | |||||||||||
| Decoder | En/Decoder | |||||||||||
| V | A | L | V | A | L | |||||||
| DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | DSC | NSD | |
| MRI-Prost. | ||||||||||||
| MRI-Heart | ||||||||||||
| MRI-Brain | ||||||||||||
| MRI-Scler. | ||||||||||||
| MRI-Hippo. | ||||||||||||
| MRI-Kidney | ||||||||||||
| MRI-Stroke. | ||||||||||||
| MRI-Resect. | ||||||||||||
| MRI-Breast | ||||||||||||
| CT-Colon | ||||||||||||
| CT-Panc. | ||||||||||||
| CT-Vessel | ||||||||||||
| US-Breast | ||||||||||||
| US-Muscle | ||||||||||||
| US-Kidney | ||||||||||||
| Xray-Hip | ||||||||||||
| Xray-Chest | ||||||||||||
| Average | ||||||||||||
| Dataset | ViT-B | MedSAM | SSLSAM | |||||||||
| Decoder | En/Decoder | Decoder | En/Decoder | Decoder | En/Decoder | |||||||
| V | A | V | A | V | A | V | A | V | A | V | A | |
| DSC | ||||||||||||
| MRI-Prost. | ||||||||||||
| MRI-Heart | ||||||||||||
| MRI-Brain | ||||||||||||
| MRI-Scler. | ||||||||||||
| MRI-Hippo. | ||||||||||||
| MRI-Kidney | ||||||||||||
| MRI-Stroke. | ||||||||||||
| MRI-Resect. | ||||||||||||
| MRI-Breast | ||||||||||||
| Average | ||||||||||||
| NSD | ||||||||||||
| MRI-Prost. | ||||||||||||
| MRI-Heart | ||||||||||||
| MRI-Brain | ||||||||||||
| MRI-Scler. | ||||||||||||
| MRI-Hippo. | ||||||||||||
| MRI-Kidney | ||||||||||||
| MRI-Stroke. | ||||||||||||
| MRI-Resect. | ||||||||||||
| MRI-Breast | ||||||||||||
| Average | ||||||||||||
| Dataset | UNet | nnUNet | SwinUNet | |||
| DSC | NSD | DSC | NSD | DSC | NSD | |
| MRI-Prost. | ||||||
| MRI-Heart | ||||||
| MRI-Brain | ||||||
| MRI-Scler. | ||||||
| MRI-Hippo. | ||||||
| MRI-Kidney | ||||||
| MRI-Stroke. | ||||||
| MRI-Resect. | ||||||
| MRI-Breast | ||||||
| Average | ||||||
| Dataset | ViT-H | ViT-B | ViT-T | |||
| H | O | H | O | H | O | |
| MRI-Prost. | ||||||
| MRI-Heart | ||||||
| MRI-Brain | ||||||
| MRI-Scler. | ||||||
| MRI-Hippo. | ||||||
| MRI-Kidney | ||||||
| MRI-Stroke. | ||||||
| MRI-Resect. | ||||||
| MRI-Breast | ||||||
| Average | ||||||
| Dataset | ViT-H | |||||
| Decoder | En/Decoder | |||||
| V | A | L | V | A | L | |
| MRI-Prost. | ||||||
| MRI-Heart | ||||||
| MRI-Brain | ||||||
| MRI-Scler. | ||||||
| MRI-Hippo. | ||||||
| MRI-Kidney | ||||||
| MRI-Stroke. | ||||||
| MRI-Resect. | ||||||
| MRI-Breast | ||||||
| Average | ||||||
| Dataset | ViT-B | |||||
| Decoder | En/Decoder | |||||
| V | A | L | V | A | L | |
| MRI-Prost. | ||||||
| MRI-Heart | ||||||
| MRI-Brain | ||||||
| MRI-Scler. | ||||||
| MRI-Hippo. | ||||||
| MRI-Kidney | ||||||
| MRI-Stroke. | ||||||
| MRI-Resect. | ||||||
| MRI-Breast | ||||||
| Average | ||||||
| Dataset | ViT-T | |||||
| Decoder | En/Decoder | |||||
| V | A | L | V | A | L | |
| MRI-Prost. | ||||||
| MRI-Heart | ||||||
| MRI-Brain | ||||||
| MRI-Scler. | ||||||
| MRI-Hippo. | ||||||
| MRI-Kidney | ||||||
| MRI-Stroke. | ||||||
| MRI-Resect. | ||||||
| MRI-Breast | ||||||
| Average | ||||||
| Dataset | ViT-H | ViT-B | ViT-T | MedSAM | SupSAM | ||||
| H | O | H | O | H | O | H | O | - | |
| MRI-Prost. | |||||||||
| MRI-Heart | |||||||||
| MRI-Brain | |||||||||
| MRI-Scler. | |||||||||
| MRI-Hippo. | |||||||||
| MRI-Kidney | |||||||||
| MRI-Stroke. | |||||||||
| MRI-Resect. | |||||||||
| MRI-Breast | |||||||||
| Average | |||||||||
| Dataset | ViT-H | |||||
| Decoder | En/Decoder | |||||
| V | A | L | V | A | L | |
| MRI-Prost. | ||||||
| MRI-Heart | ||||||
| MRI-Brain | ||||||
| MRI-Scler. | ||||||
| MRI-Hippo. | ||||||
| MRI-Kidney | ||||||
| MRI-Stroke. | ||||||
| MRI-Resect. | ||||||
| MRI-Breast | ||||||
| Average | ||||||
| Dataset | ViT-B | |||||
| Decoder | En/Decoder | |||||
| V | A | L | V | A | L | |
| MRI-Prost. | ||||||
| MRI-Heart | ||||||
| MRI-Brain | ||||||
| MRI-Scler. | ||||||
| MRI-Hippo. | ||||||
| MRI-Kidney | ||||||
| MRI-Stroke. | ||||||
| MRI-Resect. | ||||||
| MRI-Breast | ||||||
| Average | ||||||
| Dataset | ViT-T | |||||
| Decoder | En/Decoder | |||||
| V | A | L | V | A | L | |
| MRI-Prost. | ||||||
| MRI-Heart | ||||||
| MRI-Brain | ||||||
| MRI-Scler. | ||||||
| MRI-Hippo. | ||||||
| MRI-Kidney | ||||||
| MRI-Stroke. | ||||||
| MRI-Resect. | ||||||
| MRI-Breast | ||||||
| Average | ||||||